OpenStudy (anonymous):

You are walking around your neighborhood and you see a child on top of a roof of a building kick a soccer ball. The soccer ball is kicked at 33° from the edge of the building with an initial velocity of 21 m/s and lands 71 meters away from the wall. How tall is the building that the child is standing on?

OpenStudy (whpalmer4):

Okay, you need to decompose that 21 m/s @ 33 degrees into horizontal and vertical components. Do you know how to do that?

OpenStudy (anonymous):

21cos(33) ?

OpenStudy (anonymous):

and 21sin(33)?

OpenStudy (whpalmer4):

Yeah. Which is which? :-)

OpenStudy (anonymous):

21cos(33) is horizontal 31sin(33) is vertical

OpenStudy (anonymous):

answer. hor. is 17.612 vet. 11.437

OpenStudy (whpalmer4):

Okay, so we have two problems to solve: 1) what is the time of flight 2) how high is the building

OpenStudy (whpalmer4):

we can write an equation for the height of the ball as a function of time: \[h(t) = -9.8t^2 + v_0t + h_0\]where \(-9.8\) is a constant (in units of m/s^2) for the acceleration due to gravity, \(v_0\) is the initial vertical velocity (computed above), and \(h_0\) is the initial height (read: the height of the building) however, we only have 1 equation and two unknowns, so that isn't enough to find both \(t\) and \(h_0\).

OpenStudy (whpalmer4):

However, we also know that the ball travels 71 m horizontally on its trip to the ground. \[d = r*t\]\[71 m = r*t\]where \(r\) is the horizontal component of the initial velocity. With that, we can find the value of \(t\). With the value of \(t\), we can solve for \(h(t) = 0\) which will give us \(h_0\), the height of the building. Give it a shot and I'll check your answer.

OpenStudy (anonymous):

t=4.03 ho=113.069

OpenStudy (whpalmer4):

Yes, that's what I get as well.

OpenStudy (anonymous):

so the answer is 113.069?

OpenStudy (whpalmer4):

Yes, unless we both made the same mistake...

OpenStudy (whpalmer4):

Here's a plot of the height of the ball vs time (blue), the distance the ball has traveled horizontally vs. time (purple) and x = 71 meters (olive) Note that the point at which the diagonal line (the horizontal distance traveled by the ball) intersections 71 meters is exactly over the point where the height of the ball crosses the x-axis (hit the ground). That's what we want to see!